How do you write out equation when it gives you two sets of coordinates? For example:write the equation of the line:(4,-3) (5,9). What are the steps I would take to solve this problem.
Thank you for the help =](smiley)

Answer by queenofit(120)

First you would have to find your slope. this is done by using the point-slope formula. for my purposes, when you see x1 or y1 and x2 and y2, the 1's are your first set of points and the 2s are the second set of points.
the formula for finding the slope is
y2-y1
----- = m = slope for your equation
x2-x1

9-(-3)
-------= 12/1=12=m=slope
5-4
now that you have your slope you can plug it into the point slope form to get your equation that formula is y-y1=m(x-x1) for your equation the plug your numbers in
y-(-3)=12(x-4)
y+3=12x-48
y=12x-48-3
y=12x-51 This is your equation of a line in y=mx+b form